OpenStudy (cedric):

Choose the correct simplification of the expression (5x4y2z2)(3x4y3z5). 15x16y6z10 15x8y5z7 8x8y5z7 8x16y6z10

OpenStudy (australopithecus):

add the exponents of like variables and multiply the numbers

OpenStudy (cedric):

ok so hold on

OpenStudy (cedric):

is it a?

OpenStudy (australopithecus):

no

OpenStudy (australopithecus):

you add exponents when multiplying

OpenStudy (australopithecus):

x^2(x^4) = x^(2 + 4) = x^6 2 + 4 = 6

OpenStudy (anonymous):

15x8y5z7

OpenStudy (cedric):

ok so theress six but i dont get it theres no answer with x6

OpenStudy (anonymous):

5*3*x(4+4)*y(2+3)*z(2+5)

OpenStudy (cedric):

so B